Family incomes have a mean of $60 000 with a standard deviation of $20 000. The data is normally distributed. What is the probability of a randomly chosen family having an income less than $36 000? Provide evidence of your work by typing out your full solution.



1
Expert's answer
2021-08-17T10:43:34-0400

"P(X>50000) = 1-P(X\u226450000) = 1- 0.3085 = 0.6915"

Probability of a randomly chosen family having an income greater than $50 000

"P(X>50000) = 0.6915"

Probability that a family income is less than $36 000

"P(X\u226436000)"

Z-score for 36000

"=\frac{(36000-60000)}{20000 }\\
=\frac{ -24000}{20000}\\
= -1.2"

From standard normal tables , "P(Z\u2264 -1.2) =0.1151"